Chabay, Matter & Interactions, 4e Help I S Part 1 Your answer is partially correct. Try again. (a) What is the momentum of an electron traveling at a velocity of < 0, 0, -2.5x 10> m/s? 0.000000002 >kg-m/s g this Question Assistance, you will learn while you earn points based on the Point Potential Policy set b ▼ Part2 Your answer is incorrect. Try agair (b) What is the magnitude of the momentum of the electron? kg-mýs the tolerance +/-2% Assistance, you will learn while you earn points based on the Point Potential Policy set by your inst y accessing this Question you Chapter 1.5-1.13 Reading Qui
Answer #1

use the relativistic momentum formule to solve the momentum of an electron

given

m = 9.1*10^-31 kg

v = <0,0,-2.5*10^8> m/s

formule is p = me*v/\sqrt{1-v^2/c^2}

1-v^2/c^2 = 1-(-2.5*10^8)^2/(3*10^8)^2

1-v^2/c^2 = 0.3056

p = 9.1*10^-31*<0,0,-2.5*10^8>/\sqrt{}0.3056

p = <0,0,-4.11*10^-22> kg m/s

magnitude of p = 4.11*10^-22 kg m/s